OpenStudy (anonymous):

Compute derivative of f(x)=t/(2-x) and its Taylor series around a=1

OpenStudy (across):

\[f(x)=\frac{t}{2-x}=t(2-x)^{-1}\]\[f'(x)=t(2-x)^{-2}=\frac{t}{(2-x)^2}\]